OpenStudy (heart_offire):

Write the equation of the line with the given slope and y-intercept. slope = 1 - 2 y-intercept = 7 A) y = 7x + 1 - 2 B) y = 1 - 2 x + 7 C) x = 7y + 1 - 2 D) x = 1 - 2 y + 7 @pooja195

pooja195 (pooja195):

y=mx+b m=slope b=y intercept

OpenStudy (heart_offire):

so i can cross off a

pooja195 (pooja195):

m=1/2x y=7 put them into the equation :-)